Find an equation of variation if the following conditions exist.
Suppose that y varies directly as z and y = 43 when z = 387.

Answer by stanbon(48568)

Find an equation of variation if the following conditions exist.
Suppose that y varies directly as z and y = 43 when z = 387.
-----------------------------------
y = kz
--------------------
43 = k*387
k = 43/387
k = 1/9
-------------------
Equation:
y = (1/9)/z
y = 1/(9z)
=========================
Cheers,
Stan H.